Circular No. 8010

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2002 V1 S. Pravdo, Jet Propulsion Laboratory, reports the NEAT discovery, on

CCD images taken with the 1.2-m reflector at Haleakala, of a comet with a tail ∼ 10′′ long in p.a. 225^{o}. Following posting on the NEO Confirmation Page, other CCD observers have confirmed the cometary nature: P. Pravec and L. Šarounová (Ondřejov, 0.65-m reflector), coma diameter 0′.4, faint tail 0′.4 long in p.a. 250^{o}; M. Lehký (Hradec Králové, Czech Republic, 0.40-m f/5 reflector), slightly diffuse coma of diameter ∼ 5′′; M. Blasco and S. Sanchez (Mallorca, 0.40-m Schmidt telescope), coma diameter 15′′. Available astrometry:

2002 UT α_{2000} δ_{2000} m_{1} Observer
Nov. 6.59602 5^{h}05^{m}40.^{s}42 +16^{o}21′03′′.0 17.3 NEAT
6.60641 5 05 39.30 +16 21 03.9 17.5  _{”}
6.61673 5 05 38.08 +16 21 03.8 17.1  _{”}
6.82561 5 05 15.90 +16 21 19.8 16.4 Blasco
6.82802 5 05 15.80 +16 21 19.5 16.5  _{”}
6.85041 5 05 13.15 +16 21 21.9  Lehký
6.85525 5 05 12.72 +16 21 22.3  _{”}
6.85850 5 05 12.27 +16 21 21.9  _{”}
6.86370 5 05 11.67 +16 21 22.7  _{”}
6.88959 5 05 09.11 +16 21 24.7  Pravec
6.89146 5 05 08.90 +16 21 24.7  _{”}
6.89348 5 05 08.67 +16 21 24.8 16  _{”}
6.90192 5 05 07.47 +16 21 24.8  Lehký
6.90542 5 05 07.19 +16 21 26.0  _{”}
6.91061 5 05 06.62 +16 21 27.3  _{”}
6.91382 5 05 06.00 +16 21 27.1  _{”}

SUPERNOVA 2002hk IN MCG +07-15-6 A. V. Filippenko, R. Chornock, and S. Jha, University of California at

Berkeley, report that inspection of a CCD spectrum (range 400–1000 nm) obtained on Nov. 6 UT with the Keck II 10-m telescope (+ ESI) shows that SN 2002hk (cf. IAUC 8007) is a somewhat-unusual type-II supernova. Superposed on the relatively red continuum is a strong, broad Hα emission line; weaker Na D (possibly He I), Hβ with a P-Cyg profile, and several additional features are also present. The object’s spectrum bears some resemblance to that of SN 2000cs (IAUC 7448).

2002 November 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green